OpenStudy (chely2013):

f(x)=2x*e^2x. find lim of f(x) as x--> - infinity and +infinity

OpenStudy (underhill):

\[\lim_{x \rightarrow \infty}f(x)=\lim_{x \rightarrow \infty}2xe ^{2x}\]and\[\lim_{x \rightarrow -\infty}f(x)=\lim_{x \rightarrow -\infty}2xe ^{2x}\] The first limit is clearly \[+\infty\]since both 2x and e^(2x) approach infinity as x approaches infinity. The second limit is more difficult since one factor approaches negative infinity and the other approaches 0. To solve this, use L-Hospital's Rule and put e^(2x) in the denominator by making the exponent negative:\[\lim_{x \rightarrow -\infty} 2x/(e ^{-2x})\]Now we have an infinity/infinity situation to which we can apply L'Hospital's Rule. We get\[\lim_{x \rightarrow -\infty}2/(-2e^{-2x}) = \lim_{x \rightarrow -\infty}-e^{2x} = -\infty\] http://mathworld.wolfram.com/LHospitalsRule.html

OpenStudy (anonymous):

you made a mistake

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty}{2x}{e ^{2x}}\] \[\lim_{x \rightarrow -\infty}\frac {2x}{e ^{-2x}}\] \[\lim_{x \rightarrow -\infty}\frac {2}{-2e ^{-2x}}\]=0

OpenStudy (underhill):

Thanks, cinar! I stand corrected. ;-)